# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

### v1

@@ Appendix A All Prompts
Appendix B Results of Symmetry and Asymmetry
(a) Symmetry.
(b) Asymmetry.
Figure 10: Results for symmetry and asymmetry when $k=1$.
(a) Symmetry.
(b) Asymmetry.
Figure 11: Results for symmetry and asymmetry when $k=10$.
@@ Appendix C Confusion Matrices
Figure 12: Distinguishability matrices of all MLMs
Figure 13: Distinguishability matrices of all CLMs
@@ Appendix D Model Size Difference per Relation and Metric
Table 10 shows the performance difference between the largest and the smallest model for four model families. Table 11 shows the cases where a model significantly outperforms its next-larger counterpart.
Table 10: Performance difference between the largest and the smallest model for four model families. All differences in metrics where tests can be performed are significant unless they appear in a bracket. The maximum within model families and metrics is boldfaced.
Metric Relation BERT RoBERTa ALBERT OPT
Soundness HYP .02 .12 .07 .06
HPO (.00) .11 .11 .09
HOL .04 (.02) .08 .07
MER .03 .05 .12 (.01)
ANT .19 .04 .33 .25
SYN .05 .10 .14 .02
Completeness HYP .03 .04 .06 .06
HPO .05 .02 .06 .06
HOL .04 .04 .07 .06
MER .03 .04 .11 .02
ANT .16 .04 .30 .24
SYN .06 .07 .12 .03
[As/S]ymmetry HYP .04 .07 .11 .12
HPO (.00) .02 .02 .03
HOL .05 .06 .16 .10
MER .05 .05 .16 .06
ANT .15 .09 .29 .34
SYN .11 .13 .22 .04
Prototypicality HYP .03 .08 .06 .06
HOL .05 .07 .06 .05
ANT .14 (.03) .25 .22
SYN .05 .08 .12 .02
Distinguishability n.a. 1.23 1.42 4.36 4.59
Table 11: Number of smaller models who significantly outperform the next-largest model.
Metric Model Family (# Models) HYP HPO HOL MER ANT SYN
Soundness BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 0 1 0 0 0
OPT (8) 3 0 2 1 1 1
Completeness BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 1 0 0 0 0
OPT (8) 3 2 2 1 1 1
[As/S]ymmetry BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 1 0 0 1 0
OPT (8) 3 1 1 1 1 1
Prototypicality BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 n.a. 1 n.a. 0 0
OPT (8) 3 n.a. 2 n.a. 1 2
@@ Appendix F Performance Heteroscedasticity Introduced by Prompts
Table 12: Agents, given per metric and relation, present performance heteroscedasticity introduced by prompts. The column “TOTAL” shows the number of such agents out of 17 agents (16 models + humans).
@@
Soundness HYP All models 16
HPO A1, A2, A3, A4, B1, B2, R1, R2, O2, O4, O5, O6, O7, O8 14
HOL All models 16
MER A3, B2, R1, R2, O2, O3, O6 7
ANT A1, A2, A3, A4, B1, B2, O1, O2, O3, O4, O5, O6, O7, O8 14
SYN A2, A4, B2, R2, O1, O4, O5, O6, O7, O8 10
Completeness HYP All models 16
HPO A1, A2, A4, B1, B2, R2, O1, O2, O3, O4, O5, O6, O7, O8 14
HOL All models 16
MER A3, R1, R2, O2, O3, O6, O7 7
ANT A1, A2, A3, A4, B1, B2, O1, O2, O3, O4, O5, O6, O7, O8 14
SYN A4, B1, R2, O1, O3, O4, O5, O6, O7, O8 10
[As/S]ymmetry HYP All models 16
HPO All models 16
HOL ALL model 16
MER A3, B2, R1, R2, O1, O2, O3, O4, O6, O7, O8 11
ANT A1, A3, A4, O1, O2, O3, O4, O5, O6, O7, O8 11
SYN A1, A2, A4, B1, B2, R1, R2, O1, O3, O4, O5, O6, O7, O8 14

### v4

@@ Appendix A All Prompts
Appendix B Results of Symmetry
Figure 9: Results for symmetry when $k=1$.
Figure 10: Results for symmetry when $k=10$.
@@ Appendix C Confusion Matrices
Figure 11: Distinguishability matrices of all Models
@@ Appendix D Model Size Difference per Relation and Metric
Table 10 shows the performance difference between the largest and the smallest model for four model families.
Table 10: Performance difference between the largest and the smallest model for four model families. All differences in metrics are significant unless they appear in a bracket. The maximum within model families and metrics is boldfaced.
Metric Relation BERT RoBERTa Llama
Soundness HYP .02 .12 --.01
HPO (.00) .11 .02
HOL .04 (.02) (.02)
MER .03 .05 .05
ANT .19 .04 -.06
SYN .05 .10 .11
Completeness HYP .02 .04 --.01
HPO .05 .03 (.00)
HOL .04 .04 .02
MER .03 .04 .02
ANT .16 .04 --.06
SYN .06 .07 .07
Symmetry ANT .15 .09 .04
SYN .11 .13 .13
Prototypicality HYP .03 .08 --.02
HOL .05 .07 (.00)
ANT .14 (.03) --.05
SYN .05 .08 .07
@@ Appendix F Performance Heteroscedasticity Introduced by Prompts
Table 11: Agents, given per metric and relation, present performance heteroscedasticity introduced by prompts. The column “TOTAL” shows the number of such agents out of 11 agents (10 models + humans).
@@
Soundness HYP All models 6
HPO All models 6
HOL All models 6
MER B2, R1, R2, L2 4
ANT B1, B2, L1, L2 4
SYN B2, R2, L1, L2 4
Completeness HYP All models 6
HPO B1, B2, L1, L2 4
HOL All models 6
MER B2, R1, R2, L2 4
ANT B1, B2, L1, L2 4
SYN B1, R2, L1, L2 4
Symmetry ANT L1, L2 2
SYN All models 6
The prompt that achieves the best performance for each relation, metric, and agent is given in the following tables.
Table 12: Prompts leading to the best performance per relation and metric for BERT-base.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is a subordinate type of [v] 0.22
HPO [det] [w], such as [det] [v] 0.22
HOL [det] [w] is a part of [det] [v] 0.31
MER [det] [w] has [det] [v] 0.13
ANT it is not likely to be both [det] [w] and [det] [v] 0.52
SYN [det] [w] is also known as [det] [v] 0.06
Completeness HYP [det] [w] is a subordinate type of [v] 0.22
HPO [det] [w], such as [det] [v] 0.16
HOL [det] [w] is a part of [det] [v] 0.25
MER [det] [w] has [det] [v] 0.13
ANT it is not likely to be both [det] [w] and [det] [v] 0.49
SYN [det] [w] is often referred to as [det] [v] 0.13
Prototypicality HYP [det] [w] is a kind of [det] [v] 0.18
HOL [det] [w] is a part of [det] [v] 0.27
ANT it is not likely to be both [det] [w] and [det] [v] 0.41
SYN [det] [w] is often referred to as [det] [v] 0.14
Symmetry ANT it is not likely to be both [det] [w] and [det] [v] 0.61
SYN [det] [w] is often referred to as [det] [v] 0.26
Table 13: Prompts leading to the best performance per relation and metric for BERT-large.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is a subordinate type of [v] 0.34
HPO [det] [w], such as [det] [v] 0.29
HOL [det] [w] belongs to parts of [det] [v] 0.36
MER parts of [det] [w] include [det] [v] 0.22
ANT it is not likely to be both [det] [w] and [det] [v] 0.65
SYN [det] [w] is also known as [det] [v] 0.14
Completeness HYP [det] [w] is a subordinate type of [v] 0.24
HPO [det] [w], such as [det] [v] 0.21
HOL [det] [w] is a part of [det] [v] 0.26
MER parts of [det] [w] include [det] [v] 0.18
ANT it is not likely to be both [det] [w] and [det] [v] 0.62
SYN the word [w] has a similar meaning as the word [v] 0.17
Prototypicality HYP [det] [w] is a subordinate type of [v] 0.23
HOL [det] [w] is a part of [det] [v] 0.31
ANT it is impossible to be both [det] [w] and [det] [v] 0.53
SYN [det] [w] is also known as [det] [v] 0.21
Symmetry ANT it is impossible to be both [det] [w] and [det] [v] 0.73
SYN the word [w] has a similar meaning as the word [v] 0.37
Table 14: Prompts leading to the best performance per relation and metric for RoBERTa-base.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is a subordinate type of [v] 0.27
HPO [det] [w], such as [det] [v] 0.24
HOL [det] [w] is a part of [det] [v] 0.33
MER [det] [w] consists of [det] [v] 0.22
ANT [det] [w] is the opposite of [v] 0.56
SYN [det] [w] is often referred to as [det] [v] 0.13
Completeness HYP [det] [w] is a subordinate type of [v] 0.23
HPO the word [w] has a more general sense than the word [v] 0.16
HOL [det] [w] is a component of [det] [v] 0.25
MER components of [det] [w] include [det] [v] 0.18
ANT [det] [w] is the opposite of [v] 0.52
SYN the word [w] has a similar meaning as the word [v] 0.13
Prototypicality HYP [det] [w] is a subordinate type of [v] 0.22
HOL [det] [w] is a part of [det] [v] 0.28
ANT [det] [w] is the opposite of [v] 0.50
SYN [det] [w] is also called [det] [v] 0.18
Symmetry ANT [det] [w] is the opposite of [v] 0.59
SYN [det] [w] is often referred to as [det] [v] 0.17
Table 15: Prompts leading to the best performance per relation and metric for RoBERTa-large.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is a type of [v] 0.47
HPO [det] [w], such as [det] [v] 0.39
HOL [det] [w] is a part of [det] [v] 0.33
MER parts of [det] [w] include [det] [v] 0.33
ANT if something is [det] [w], then it can not also be [det] [v] 0.60
SYN [det] [w] is also called [det] [v] 0.27
Completeness HYP [det] [w] is a type of [v] 0.26
HPO the word [w] has a more general sense than the word [v] 0.20
HOL [det] [w] is a part of [det] [v] 0.28
MER components of [det] [w] include [det] [v] 0.26
ANT if something is [det] [w], then it can not also be [det] [v] 0.58
SYN [det] [w] is also called [det] [v] 0.21
Prototypicality HYP [det] [w] is a type of [v] 0.34
HOL [det] [w] is a part of [det] [v] 0.36
ANT if something is [det] [w], then it can not also be [det] [v] 0.52
SYN [det] [w] is also known as [det] [v] 0.28
Symmetry ANT [det] [w] is the opposite of [v] 0.73
SYN [det] [w] is also known as [det] [v] 0.34
Table 16: Prompts leading to the best performance per relation and metric for LLaMA-3-8B.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is [det] [v] 0.42
HPO the word [w] has a more general sense than the word [v] 0.30
HOL [det] [w] is a constituent of [det] [v] 0.39
MER constituents of [det] [w] include [det] [v] 0.08
ANT the word [w] has an opposite meaning of the word [v] 0.86
SYN the word [w] means nearly the same as the word [v] 0.55
Completeness HYP [det] [w] is a kind of [det] [v] 0.21
HPO the word [w] has a more general sense than the word [v] 0.14
HOL [det] [w] is a part of [det] [v] 0.27
MER components of [det] [w] include [det] [v] 0.08
ANT the word [w] has an opposite meaning of the word [v] 0.82
SYN the word [w] means nearly the same as the word [v] 0.38
Prototypicality HYP [det] [w] is a kind of [det] [v] 0.29
HOL [det] [w] is a constituent of [det] [v] 0.32
ANT the word [w] has an opposite meaning of the word [v] 0.68
SYN the word [w] means nearly the same as the word [v] 0.39
Symmetry ANT the word [w] has an opposite meaning of the word [v] 0.88
SYN the word [w] has a similar meaning as the word [v] 0.50
Table 17: Prompts leading to the best performance per relation and metric for LLaMA-3-70B.
Metric Relation Prompt Performance
Soundness HYP [det] [w] is a kind of [det] [v] 0.46
HPO the word [w] has a more general meaning than the word [v] 0.31
HOL [det] [w] belongs to parts of [det] [v] 0.45
MER constituents of [det] [w] include [det] [v] 0.19
ANT the word [w] has an opposite meaning of the word [v] 0.84
SYN the word [w] has a similar meaning as the word [v] 0.65
Completeness HYP [det] [w] is a kind of [det] [v] 0.22
HPO the word [w] has a more general sense than the word [v] 0.15
HOL [det] [w] is a part of [det] [v] 0.34
MER constituents of [det] [w] include [det] [v] 0.14
ANT the word [w] has an opposite meaning of the word [v] 0.81
SYN the word [w] means nearly the same as the word [v] 0.44
Prototypicality HYP [det] [w] is a kind of [det] [v] 0.31
HOL [det] [w] belongs to parts of [det] [v] 0.34
ANT the word [w] has an opposite meaning of the word [v] 0.67
SYN the word [w] has a similar meaning as the word [v] 0.44
Symmetry ANT the word [w] has an opposite meaning of the word [v] 0.92
SYN the word [w] has a similar meaning as the word [v] 0.66